40 liters of milk contain 10% water. find the quantity of water added to the milk so that it may contain 20% water
Answer by mananth(16949)

--------percent ---------------- quantity
Milk 10 ---------------- 40 L
Water 100 ---------------- x L
Mixture 20 ---------------- 40 + x L

10*40+100x=20(40+x)
400+100 x=800+20x
100x-20 x=800+400
80x=400
/ 80
x= 5
5L of Water